# Behavioral and Topological Heterogeneities in Network Versions of Schelling’s Segregation Model

### v3

@@ 5.3   Heterogeneous Topology
This method ensures that neighborhood densities are distributed consistently across the network, so nodes and their neighbors cannot have unrealistic differences in degree. The result of multiple densifications is substantially greater variation in node degree across the network as well as a marked increase in the mean node degree. The result is a variety of neighborhood sizes. While in the base model, each neighborhood is a von Neumann neighborhood bordering another von Neumann neighborhood (except at the boundary), randomly densified lattices have a variety of neighborhood boundary relationships, e.g., a von Neumann neighborhood where individuals have 4 neighbors, might be adjacent to a Moore neighborhood where individuals have 8 neighbors. This enables a richer diversity of neighborhood relationships.
@@
It is clear that the densification procedure effectively produces the desired heterogeneity of topology. The level of topological heterogeneity produced at 128 densifications exceeds the level produced by any network generator in [4]. Because there is an upper bound on local density, successive increases in densifications lead to progressively smaller increases in heterogeneity of neighborhood size and neighborhood-size pair configurations. Eventually, continued densification returns the topology to homogeneity.
@@ 6   Experiments & Results
For the primary set of experiments, ten unique parameter configurations were used: Half employed uniform Schelling [2] tolerance threshold assignments and half used heterogenous Xie & Zhou [3] tolerance threshold assignments. For each tolerance assignment method, simulations were conducted with 0, 32, 64, 96, and 128 densifications. For each parameter setting, 100 simulation runs were conducted. For each simulation, 4000 timesteps were monitored. To ensure robustness of results, additional simulations were conducted with increased vacancy on the lattice and with the presence of a minority population.
@@ 6.1   Segregation
For convenience, Newman’s [28] assortativity coefficient, $r$, is used as a measure of segregation levels during the simulations. $r$ is given by Eq. 2 in [28]:
@@
$\frac{\sum_{i}e_{ii}-\sum_{i}{a_{i}b_{i}}}{1-\sum_{i}{a_{i}b_{i}}}\approx\frac{1-e_{ij}-.5}{1-.5}=1-2e_{ij}$ (5)
Values of $r$ close to 1 indicate high levels of segregation while values of $r$ close to 0 indicate approximately random mixing. To account for vacant nodes on the network, the assortativity coefficient for the subgraph containing only occupied nodes is used. For all simulations where the derived subgraph consisted of multiple connected components, $r$ was calculated for each connected component and a weighted average was constructed:
@@
where $N$ is the total number of occupied nodes and $r_{i}$ and $n_{i}$ are the assortativity coefficient and size of the $i^{th}$ component, respectively. For connected components with homogenous composition, a value of 1 was assigned for $r$. To observe the process of segregation, $r^{\prime}$ was recorded at each time step.
The observed mean final assortativity was higher for all Schelling-tolerance simulation runs. As in Xie & Zhou [3], heterogeneity of tolerances did, on their own, result in reduced assortativity. Consistent with Gandica, Gargiulo, & Carletti’s [4] observations, increasing topological heterogeneity alone did not produce results with reduced assortativity when $\epsilon\approx 0.39$, in fact, any increase in topological heterogeneity for Schelling-tolerance simulations resulted in a marked increase in realized assortativity. The combination of both dimensions of heterogeneity resulted in progressive reductions in assortativity. Xie & Zhou-tolerance simulations generated assortativity more slowly and resulted in reduced levels of segregation compared with Schelling-tolerance simulations at every level of densification. The lowest levels of assortativity observed occurred with Xie & Zhou [3] tolerances and 128 densifications. Results are illustrated in Figure 4.
@@ 6.2.1   Hub Composition
The mean degree for each graph and its vacancies was calculated; see Figure 5. The mean vacancy degree showed greater variability than the graph degree for each group. No significant difference was observed between the accumulations of vacancies in Xie & Zhou tolerance simulations and Schelling tolerance simulations. When substantial levels of densification were present, the mean vacancy degree could exceed the mean graph degree.
@@
Otherwise, the mean vacancy degree tended lower. For simulations with Xie & Zhou tolerances on densified graphs, on average, individuals in the group with the highest tolerance thresholds ($\geq 0.57$) had distinctly higher degree centralities than those in the group with the lowest tolerance thresholds ($\leq 0.07$). This difference became more pronounced as the number of densifications increased. These results are illustrated in Figures 6-7. These results were robust to increases in excess housing and the presence of a minority population.
Figure 6: Final mean degree by node tolerance and level of densification.
@@ 6.2.2   Paired Tolerances
In addition to the topological organization of tolerances, the pairwise organization of tolerances can also be observed. At every level of densification, the distribution of pairs’ tolerances became organized as segregation increased. This effect became less pronounced as the number of densifications increased. Additionally, the organization of tolerances can be observed visually in Figure 7. At all levels of densification, distinct regions of like-tolerance nodes can be observed. This organization is further illustrated in Figure 8 where the relative frequencies of tolerance-pairs for dissimilar neighbors are shown. The overall level of organization exhibited at each level, the difference in levels of heterogeneity, is shown in Table 1. As densifications increase, final graphs are less restrictively organized.
Figure 8: Heatmaps show the log frequencies of tolerance-level pairs for dissimilar neighbors. Left, 0 densifications; right, 128 densifications.
@@
Figure 9: This heatmap shows the log frequencies of tolerance-level pairs for dissimilar neighbors when the excess housing rate is 30%.
@@ 7.1   Ordered Migration
where $\epsilon$ is the individual’s tolerance threshold and N is the neighborhood size. Since our models begin with a randomly mixed population where $r\approx 0$ and $r$ increases monotonically, we need only consider cases where $0\leq p\leq 0.5$, as shown in Figure10.For $\epsilon=0.39$, with only heterogeneity of topology, individuals in smaller neighborhoods will generally become more likely to exceed their tolerance thresholds than those in larger neighborhoods as assortativity increases over time. In this way, an additional dynamic is included: the out-migration of non-dominant-type individuals in densified clusters will precede in-migration to those clusters by dominant-type individuals. This behavior enables the increase in assortativity observed in Schelling-tolerance simulations with topological heterogeneity. When both topological and behavioral heterogeneities are represented, dynamics are richer. The probability of an individual’s residence in a location that exceeds its tolerance threshold can be roughly ordered from greatest to least: low tolerance, large neighborhood; low tolerance, small neighborhood; moderate tolerance, small neighborhood; moderate tolerance, large neighborhood; high tolerance, small neighborhood; and high tolerance, large neighborhood.
Figure 10: Vertical axis: Probability of exceeding tolerance threshold, $P_{\epsilon}$. Horizonal axis: probability of $n^{th}$ dissimilar neighbor, $p$. Note that as assortativity increases, the probability of dissimilar neighbors decreases.
@@ 7.2   Tolerance Repels Intolerance
When heterogeneity of tolerances is represented, pairs of dissimilar individuals with high tolerance thresholds create another interesting dynamic: these nodes repel intolerant nodes and attract tolerant neighbors. This dynamic is restrained on a 2D grid with von Neumann neighborhoods. Such pairs can have no common neighbors, but some of the neighbors of each will share an edge; see Figure 11. With network topologies where dissimilar adjacent nodes may share a neighbor, that neighbor must possess sufficient tolerances to accommodate at least one pair of dissimilar neighbors.
@@
When both heterogeneous tolerances and heterogeneous topology are represented, the influence of tolerant nodes is amplified by densifications. A densified cluster populated by dissimilar, tolerant nodes acts as a strong repellant for intolerant nodes as tolerant pairs will directly share more neighbors. Second, a densified cluster occupied by similar nodes may continue to attract similar nodes but can no longer repel a dissimilar individual with a sufficiently large tolerance threshold. The introduction of relatively few highly tolerant dissimilar individuals to such a cluster will motivate catastrophic outmigration of intolerant individuals. These dynamics lead some densified clusters to become stores of tolerance and diversity. As the level of densification increases, the presence of tolerant nodes and densified clusters makes extremely homogeneous areas rarer. As this occurs, the most intolerant nodes can no longer find suitably homogenous candidate vacancies and increasingly remain dissatisfied and connected to tolerant, dissimilar nodes; see Figure 8. Thus, both the organization of tolerances and the organization of color-types in the final networks are reduced; see Table 1. It is important to note, however, that the presence of excess housing can moderate these dynamics by providing intolerant individuals with more opportunities to avoid dissimilar neighbors.
@@ 8   Conclusion and Future Work
In this study, we investigated the impact of representing multiple dimensions of heterogeneity in network models of residential segregation. Models which combine heterogeneity of tolerances with heterogeneity of topologies showed substantially different behavior than those which employed only one dimension of heterogeneity. Ordered migration and the resultant clusters of tolerance break down intolerant homogeny. This has bearing on the important question raised previously: to what extent can the collective dynamics of individual preferences lead to residential segregation? These results indicate that these collective dynamics may not contribute to residential segregation as much as previously thought if sufficient heterogeneity is represented. At a minimum, these results indicate the necessity of representing both heterogeneity of tolerances and heterogeneity of topologies in network models of residential segregation, as the omission of one or the other will result in the loss of these important model behaviors.
@@
This work has implications for social and organizational cultural diffusion. The results suggest that the organization of tolerance levels in a network can be influenced by the network’s topology. As such, the organization of cultural diffusion in a network may be influenced by the network’s topology. This effect would be most pronounced in social and organizations systems where segregation along cultural lines is likely to occur. Future work should investigate the impact of network topology on cultural diffusion in social and organizational systems.
In addition, this work may hold importance for policymakers. The results suggest that heterogeneity of topology can either moderate or amplify segregation depending on the proportion and density of available excess housing. If policymakers aim to reduce segregation in communities, they may consider replacing low-density housing with high-density housing. Simply continuing to add both types of housing stock will continue to provide opportunities for tolerant individuals to self-organize in high-density areas while intolerant individuals continue to retreat to low-density areas. Given the variability and flexibility of zoning ordinances in the United States, this may never be a feasible policy option. A comparative analysis of real-world mixing patterns in high-density and low-density areas would be a valuable next step in this research. Another approach may examine the proportions of low- and high-density housing included in municipalities’ development plans. Emphasis on low-density housing may be a predicator of future segregation and urban-rural cultural fragmentation while underemphasis on low-density housing may predict a decline in cultural diversity.

### v4

@@ 4.2   Heterogeneous Topology
This method ensures that neighborhood densities are distributed consistently across the network, so nodes and their neighbors cannot have unrealistic differences in degree. The result of multiple densifications is substantially greater variation in node degree across the network as well as a marked increase in the mean node degree. The result is a variety of neighborhood sizes. While in the base model, all neighborhoods are von Neumann neighborhoods (except at the boundary), randomly densified lattices have a variety of neighborhood boundary relationships, e.g., a von Neumann neighborhood where individuals have 4 neighbors, might be adjacent to a Moore neighborhood where individuals have 8 neighbors. This enables a richer diversity of neighborhood relationships.
@@
It is clear that the densification procedure effectively produces the desired heterogeneity of topology. The level of topological heterogeneity produced at 128 densifications exceeds the level produced by any network generator in [4]. Because there is an upper bound on local density, successive increases in densifications lead to progressively smaller increases in heterogeneity of neighborhood size and neighborhood-size pair configurations. Eventually, continued densification returns the topology to homogeneity. Here, we ensure our densification procedure does not inadvertently produce a homogeneous topology by limiting the number of densifications to 128.
@@ D $\mu$ $\sigma$ $\eta$ Hk Hpairs $H(T)_{0}$ $H(T)_{F}$
4.3   Heterogeneous Tolerances
Following Xie & Zhou [3], agents are provided with heterogeneous tolerances (tolerance thresholds) aligned with the Guttman scale and rank-ordered logit model derived from Bruch & Mare’s [5] Detroit data. Tolerance thresholds were assigned by drawing values from a uniform distribution over a given interval: For 10.47% of individuals, $\epsilon_{i}$ fell within $[0.0,0.07)$; for 18.10% of individuals, $\epsilon_{i}$ fell within $[0.07,0.21)$; for 26.73% of individuals, $\epsilon_{i}$ fell within $[0.21,0.36)$; for 13.86% of individuals $\epsilon_{i}$ fell within $[0.36,0.57)$; for 26.59% of individuals, $\epsilon_{i}$ fell within $[0.57,1.00]$. For these individuals, the simulation procedure described in the previous section was implemented with $\epsilon_{i}$ replacing $\epsilon$. Figure 3 shows an example distribution of tolerance thresholds drawn from the constructed distribution. For the 4.25% of individuals in the Detroit data found not to conform to the Guttman scale, no static tolerance threshold was set. Instead, Xie & Zhou’s rank-ordered logit model (Eq. 4 in [3]) was implemented to determine the probability of transition to each candidate neighborhood. The transition destination, $j\in C_{Vt}$, is then randomly selected with probability, $\widehat{p_{jt}}$, given by the model:
[…]
The adoption of Xie & Zhou’s tolerance model increases the Shannon entropy of tolerance thresholds, $H_{\epsilon}$, from 00 in the homogeneous Schelling case to at least $H_{\epsilon}\approx 5.102$. This approximation was calculated by discretizing assigned tolerance thresholds into 37 identically sized bins and allowing non-Guttman scale agents to occupy an additional distinct bin.
@@ 5   Experiments & Results
For the primary set of experiments, ten unique parameter configurations were used: Half employed uniform Schelling [2] tolerance threshold assignments and half used heterogenous Xie & Zhou [3] tolerance threshold assignments. For each tolerance assignment method, simulations were implemented in Python and conducted with 0, 32, 64, 96, and 128 densifications. For each parameter setting, 100 simulations were run for 4000 time steps. Convergence was defined as the cumulative change in assortativity remaining below 0.01 over any 100-step rolling window. All simulations met this criterion within the allotted timeframe. To ensure robustness of results, additional simulations were conducted with increased vacancy on the lattice and with the presence of a minority population.
@@ 5.1   Segregation
Newman’s [38] assortativity coefficient, $r$, a well-established measure of categorical mixing on networks, is used as a measure of segregation levels during the simulations. $r$ is given by Eq. 2 in [38]:
@@
$r=\frac{\sum_{i}e_{ii}-\sum_{i}{a_{i}b_{i}}}{1-\sum_{i}{a_{i}b_{i}}}\approx\frac{1-e_{ij}-.5}{1-.5}=1-2e_{ij}$ (5)
[…]
@@
where $N$ is the total number of occupied nodes and $r_{i}$ and $n_{i}$ are the assortativity coefficient and size of the $i^{th}$ component, respectively. For connected components with homogeneous composition, a value of 1 was assigned for $r$. To observe the process of segregation, $r^{\prime}$ was recorded at each time step.
The observed mean final assortativity was higher for all Schelling-tolerance simulation runs. As in Xie & Zhou [3], heterogeneity of tolerances did, on their own, result in reduced assortativity. Consistent with Gandica, Gargiulo, & Carletti’s [4] observations, increasing topological heterogeneity alone did not produce results with reduced assortativity when $\epsilon\approx 0.39$, in fact, any increase in topological heterogeneity for Schelling-tolerance simulations resulted in a marked increase in realized assortativity. The combination of both dimensions of heterogeneity resulted in progressive reductions in assortativity. Xie & Zhou-tolerance simulations generated assortativity more slowly and resulted in reduced levels of segregation compared with Schelling-tolerance simulations at every level of densification. The lowest levels of assortativity observed occurred with Xie & Zhou [3] tolerances and 128 densifications. Results are illustrated in Figures 4 and 5.
Figure 4: Average assortativity over time for each group of simulations.
@@ 5.2.1   Hub Composition
The mean degree for each graph and its vacancies was calculated; see Figure 6. The mean vacancy degree showed greater variability than the graph degree for each group. No significant difference was observed between the accumulations of vacancies in Xie & Zhou tolerance simulations and Schelling tolerance simulations. When substantial levels of densification were present, the mean vacancy degree could exceed the mean graph degree. Otherwise, the mean vacancy degree tended lower.
@@
[…]
Densification Comparison $D$-statistic $p$-value Centrality with max diff
0 Intolerant vs Moderate 0.053 $1.58\times 10^{-19}$ 3
Intolerant vs Tolerant 0.063 $1.08\times 10^{-22}$ 3
Moderate vs Tolerant 0.009 0.119 3
32 Intolerant vs Moderate 0.274 $<10^{-300}$ 4
Intolerant vs Tolerant 0.163 $3.32\times 10^{-152}$ 8
Moderate vs Tolerant 0.115 $6.65\times 10^{-182}$ 4
64 Intolerant vs Moderate 0.382 $<10^{-300}$ 8
Intolerant vs Tolerant 0.295 $<10^{-300}$ 8
Moderate vs Tolerant 0.094 $4.32\times 10^{-122}$ 4
96 Intolerant vs Moderate 0.395 $<10^{-300}$ 8
Intolerant vs Tolerant 0.329 $<10^{-300}$ 11
Moderate vs Tolerant 0.073 $5.98\times 10^{-74}$ 4
128 Intolerant vs Moderate 0.379 $<10^{-300}$ 11
Intolerant vs Tolerant 0.358 $<10^{-300}$ 12
Moderate vs Tolerant 0.041 $3.77\times 10^{-23}$ 4
Table 2: K-S Test Results by Densification Level: Centrality Differences Among Tolerance Groups
Figure 7: Empirical cumulative distribution functions (ECDFs) of degree centrality for tolerant, moderate, and intolerant individuals across different densification levels. The plots illustrate the increasing divergence in centrality distributions as network density increases.
@@ 5.2.2   Paired Tolerances
In addition to the topological organization of tolerances, the pairwise organization of tolerances can also be observed. At every level of densification, the distribution of pairs’ tolerances became organized as segregation increased. This effect became less pronounced as the number of densifications increased. At all levels of densification, distinct regions of like-tolerance nodes can be observed. This organization is further illustrated in Figure 9 where the relative frequencies of tolerance-pairs for dissimilar neighbors are shown. The overall level of organization exhibited at each level, the difference in levels of heterogeneity, is shown in Table 1. As densifications increase, final graphs are less organized.
Figure 9: Heatmaps show the log frequencies of tolerance-level pairs for dissimilar neighbors. Left, 0 densifications; right, 128 densifications. A tolerance level of -1 denotes non-Guttman individuals.
@@
Figure 10: This heatmap shows the log frequencies of tolerance-level pairs for dissimilar neighbors when the excess housing rate is 30%. A tolerance level of -1 denotes non-Guttman individuals.
@@ 6   Discussion
The results of this study show that the introduction of multiple dimensions of heterogeneity in network models of residential segregation can lead to significant changes in the dynamics of segregation. The combination of topological and behavioral heterogeneities produces reductions in segregation beyond those observed with either dimension alone. In addition, the presence of these complementary heterogeneities lead to the emergence of two notable dynamics: ordered migration and tolerance repelling intolerance.
@@ 6.1   Ordered Migration
where $\epsilon$ is the individual’s tolerance threshold and N is the neighborhood size. Since our models begin with a randomly mixed population where $r\approx 0$ and $r$ increases approximately monotonically, we need only consider cases where $0\leq p\leq 0.5$, as shown in Figure 11. As the migration process unfolds, $p$ decreases. For $\epsilon=0.39$, with only heterogeneity of topology, individuals in smaller neighborhoods will generally become more likely to exceed their tolerance thresholds than those in larger neighborhoods as assortativity increases over time. In this way, an additional dynamic is included: the out-migration of non-dominant-type individuals in densified clusters will precede in-migration to those clusters by dominant-type individuals. This behavior enables the increase in assortativity observed in Schelling-tolerance simulations with topological heterogeneity. When both topological and behavioral heterogeneities are represented, dynamics are richer. The probability of an individual’s residence in a location that exceeds its tolerance threshold can be roughly ordered from greatest to least: low tolerance, large neighborhood; low tolerance, small neighborhood; moderate tolerance, small neighborhood; moderate tolerance, large neighborhood; high tolerance, small neighborhood; and high tolerance, large neighborhood. This ordering holds for $0.1\leq p\leq 0.4$. While this model assumes independent and identically distributed neighbor types—an assumption that holds approximately only at initialization—its predictions serve as a theoretical baseline, offering valid insight into early-stage dynamics and comparative sensitivity to dissimilarity under varying tolerance thresholds and neighborhood sizes.
Figure 11: Vertical axis: Probability of exceeding tolerance threshold, $P_{\epsilon}$. Horizontal axis: probability of $n^{th}$ dissimilar neighbor, $p$. Note that as assortativity increases, the probability of dissimilar neighbors decreases.
@@ 6.2   Tolerance Repels Intolerance
When heterogeneity of tolerances is represented, pairs of dissimilar individuals with high tolerance thresholds create another interesting dynamic: these nodes repel intolerant nodes and attract tolerant neighbors. This dynamic is restrained on a 2D grid with von Neumann neighborhoods. Such pairs can have no common neighbors, but some neighbors of each will share an edge; see Figure 12. With network topologies where dissimilar adjacent nodes may share a neighbor, that neighbor must possess sufficient tolerances to accommodate at least one pair of dissimilar neighbors.
@@
When both heterogeneous tolerances and heterogeneous topology are represented, the influence of tolerant nodes is amplified by densifications. A densified cluster populated by dissimilar, tolerant nodes acts as a strong repellant for intolerant nodes as tolerant pairs will directly share more neighbors. Second, a densified cluster occupied by similar nodes may continue to attract similar nodes but can no longer repel a dissimilar individual with a sufficiently large tolerance threshold. The introduction of relatively few highly tolerant dissimilar individuals to such a cluster will motivate catastrophic out-migration of intolerant individuals. These dynamics lead some densified clusters to become stores of tolerance and diversity. As the level of densification increases, the presence of tolerant nodes and densified clusters makes extremely homogeneous areas rarer. As this occurs, the most intolerant nodes can no longer find suitably homogeneous candidate vacancies and increasingly remain dissatisfied and connected to tolerant, dissimilar nodes; see Figure 9. Thus, both the organization of tolerances and the organization of color-types in the final networks are reduced; see Table 1. It is important to note, however, that the presence of excess housing can moderate these dynamics by providing intolerant individuals with more opportunities to avoid dissimilar neighbors.
Our models did not incorporate random movements by otherwise satisfied individuals. This likely moderated tolerant individuals’ ability to monopolize densified clusters. The presence of such random movements would likely lead to a more pronounced effect of tolerance repelling intolerance. This is a critical avenue for future work. Rather, to the extent that it is observed, tolerant individuals’ dominance of densified clustered is an example of sensitivity to initial conditions. The aforementioned entrenchment of tolerant individuals in densified clusters at initialization is, on its own, sufficient to produce this dynamic.
@@ 7   Conclusion and Future Work
In this study, we investigated the impact of representing multiple dimensions of heterogeneity in network models of residential segregation. Models which combine heterogeneity of tolerances with heterogeneity of topologies showed substantially different behavior than those which employed only one dimension of heterogeneity. Ordered migration and the resultant clusters of tolerance break down intolerant homogeny. This has bearing on the important question raised previously: to what extent can the collective dynamics of individual preferences lead to residential segregation? These results indicate that these collective dynamics may not contribute to residential segregation as much as previously thought if sufficient heterogeneity is represented. At a minimum, these results indicate the necessity of representing both heterogeneity of tolerances and heterogeneity of topologies in network models of residential segregation, as the omission of one or the other will result in the loss of these significant model behaviors.
@@
This work has implications for social and organizational cultural diffusion. The results suggest that the organization of tolerance levels in a network can be influenced by the network’s topology. As such, the organization of cultural diffusion in a network may be influenced by the network’s topology. This effect would be most pronounced in social and organizations systems where segregation along cultural lines is likely to occur. Future work should investigate the impact of network topology on cultural diffusion in social and organizational systems. Additionally, this work did not consider the effect of economic or other social factors on the dynamics of segregation. In Zhang, a notable result arising from asymmetric preferences was that individual who prefer homogeneity seemed to congregate [10]. Our result is consistent with this observation, however the mechanism is different, and the congregation is observed in sparse areas. Where Zhang employed simulated economic incentives, our result shows that topological heterogeneity can motivate aggregation of intolerant individuals as well. Future work incorporating these factors may provide additional insight into the dynamics of segregation in networks with multiple dimensions of heterogeneity.
This research may hold importance for policymakers. The results suggest that heterogeneity of topology can either moderate or amplify segregation depending on the proportion and density of available excess housing. If policymakers aim to reduce segregation in communities, they may consider replacing low-density housing with high-density housing. Simply continuing to add both types of housing stock will continue to provide opportunities for tolerant individuals to self-organize in high-density areas while intolerant individuals continue to retreat to low-density areas. Given the variability and flexibility of zoning ordinances in the United States, this may never be a feasible policy option. A comparative analysis of real-world mixing patterns in high-density and low-density areas would be a valuable next step in this research. Another approach may examine the proportions of low- and high-density housing included in municipalities’ development plans. Emphasis on low-density housing may be a predictor of future segregation and urban-rural cultural fragmentation while underemphasis on low-density housing may predict a decline in cultural diversity.
8   Acknowledgement
A Thesis has previously been published [40].